OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!!

OpenStudy (abbycross167):

Please help me to simplify- 4 ------- 12x+16

OpenStudy (rebecca):

Simplify. All terms can be divided by 4. So divide 4 from all.

OpenStudy (abbycross167):

so 4/12=3 4/16 = 4 then what @Rebecca

OpenStudy (thatonegirl_):

well with the denominator there is a common factor of 4 correct?

OpenStudy (thatonegirl_):

You can divide each term by 4 and get 4(3x+4) for the denominator

OpenStudy (thatonegirl_):

and the numerator's 4 and the one from the denominator will cancel out and you are left with just \[\frac{ 1 }{ 3x+4 }\]

OpenStudy (mathmale):

Abby, any further questions regarding this problem?